OpenStudy (anonymous):

Jyoti and Sushma together can do a piece of work in 6 days and Sushma alone can do it un 9 days .In how many days can Jyoti alone complete the work ?

OpenStudy (anonymous):

In my first thought, I would look at it like this, J + S = 6. S = 9...what is J

OpenStudy (anonymous):

Let the total work 18k 9J=18k For S 18k/9=2k per day S=2k 6J+6S=18k(for working 6 days) 6J+6(2k)=18k J=k per day so for J doing 18k alone 18k/k=18 days

OpenStudy (anonymous):

can u explain the solution ?

OpenStudy (anonymous):

I give the total work 18k then i write 2 equations 1) 6j+6s=18k these 6's are the days they work together 2) 9s=18k 9 is the day s work alone now for 2. equation s=2k then put s=2k in first equation that gives 6j+6(2k)=18k 6j+12k=18k 6j=6k so j=k j can work k per day we have 18k work and k per day worker just divide them 18k/k=18days

OpenStudy (anonymous):

but why the total work as "18k'

OpenStudy (anonymous):

i take it for easy calculations .6 and 9 can divide 18 right if u need something like that just multiply both numbers and use it the answer will be the same like 6.9=54k

OpenStudy (anonymous):

OHH u hv taken the Least Common Multiple.......

OpenStudy (anonymous):

yea. just write equations for this kind of porblems

OpenStudy (anonymous):

6J+6(2k)=18k(WHAT DO U MEAN BY THAT ?)

OpenStudy (anonymous):

s=2k right. i put s in the equation

OpenStudy (anonymous):

WHAT IS K ?

OpenStudy (anonymous):

lol sorry k is just a unit that i use in math equations

OpenStudy (anonymous):

THANKS